# footnotetext: * Equal contribution, Corresponding author

### v1

@@ Abstract
Securing personal identity against deepfake attacks is increasingly critical in the digital age, especially for celebrities and political figures whose faces are easily accessible and frequently targeted. Most existing deepfake detection methods focus on general-purpose scenarios and often ignore the valuable prior knowledge of known facial identities, e.g., "VIP individuals" whose authentic facial data are already available. In this paper, we propose VIPGuard, a unified multimodal framework designed to capture fine-grained and comprehensive facial representations of a given identity, compare them against potentially fake or similar-looking faces, and reason over these comparisons to make accurate and explainable predictions. Specifically, our framework consists of three main stages. First, fine-tune a multimodal large language model (MLLM) to learn detailed and structural facial attributes. Second, we perform identity-level discriminative learning to enable the model to distinguish subtle differences between highly similar faces, including real and fake variations. Finally, we introduce user-specific customization, where we model the unique characteristics of the target face identity and perform semantic reasoning via MLLM to enable personalized and explainable deepfake detection. Our framework shows clear advantages over previous detection works, where traditional detectors mainly rely on low-level visual cues and provide no human-understandable explanations, while other MLLM-based models often lack a detailed understanding of specific face identities. To facilitate the evaluation of our method, we built a comprehensive identity-aware benchmark called VIPBench for personalized deepfake detection, involving the latest 7 face-swapping and 7 entire face synthesis techniques for generation. Extensive experiments show that our model outperforms existing methods in both detection and explanation. All code and datasets will be released upon acceptance.
@@ 1 Introduction
The rapid advancement of generative AI techniques [46, 54, 63, 68, 47, 12] has led to the widespread creation and dissemination of deepfake content—synthetic media where a person’s identity is manipulated or replaced, often without consent. These techniques severely threaten personal reputation and public trust, especially for high-profile individuals such as celebrities, political figures, and public officials [2]. To this end, protecting one’s identity from such manipulations has become more than just a personal privacy issue—it is a pressing societal concern.
@@
Although deepfake detection has gained attention, most existing methods are designed for general-purpose use [45, 33, 51, 69, 65, 41]. They aim to classify any face image or video as real or fake, without considering whose face is being targeted. In the real world, many scenarios provide access to prior knowledge about the target identity [^1] In this work, we refer to such individuals as ”VIPs” whose identities we aim to protect., such as in the case of public figures or known individuals [15]. This opens a new avenue for detection: Can we leverage the known facial identity to improve both detection and explainability in personalized deepfake detection? Rather than treating all faces the same, identity-aware detection methods [49, 18, 16, 15, 73] focus on the semantic alignment between the input image and the authentic identity. By doing so, they provide more personalized and context-aware predictions, potentially improving both detection performance and explainability.
@@
However, existing identity-aware detectors [15, 73] fail to fully utilize the detailed identity-specific information. They primarily rely on global facial features while neglecting fine-grained semantic details—such as eye shape, facial contours, or other attributes. As shown in Figure 1, an image generated by the latest GPT-4o [43] can appear highly realistic at first glance, but still contains subtle inconsistencies in local facial regions—for example, unusually pronounced eye bags. When the target identity is known and all facial details are available, leveraging these fine-grained discrepancies for detection becomes especially promising.
@@
Figure 1: Comparison between natural images and images generated by GPT-4o [43]. Inconsistencies can be observed in localized facial attributes, such as the eye pouches and facial shapes.
@@
To this end, in this paper, we propose VIPGuard, a unified multimodal framework for detecting and explaining deepfakes targeting specific users. VIPGuard addresses identity-aware deepfake detection by explicitly incorporating known facial identity priors, including both global identity information and detailed structural facial attributes from the VIPs. To achieve this, this paper, for the first time, reformulates forgery detection as a fine-grained face recognition task, targeted at VIP identities. The framework consists of three stages. To leverage both global and local facial information, we use pre-trained face models [57, 13, 10] to extract global facial priors (face similarity scores) and local facial priors (facial attributes). Leveraging these priors, we aim to enable an MLLM to perform forgery detection through semantic comparisons between suspect and authentic faces across visual and textual inputs. We train VIPGuard in three main stages: (1) Fine-tuning an MLLM with a variety of facial attribute data to enhance facial understanding; (2) Performing discriminative learning to distinguish some arbitrary identities from manipulated or similar-looking faces by reasoning; (3) Supporting personalized detection by learning a unique and lightweight VIP token to represent each target identity for customized reasoning.
@@
To enable robust evaluation in personalized deepfake detection, we additionally present VIPBench—a comprehensive and identity-centric benchmark that differs fundamentally from conventional deepfake benchmarks [70, 67]. While existing benchmarks typically treat faces generically and overlook whose identity is being manipulated, VIPBench focuses explicitly on identity-aware scenarios, where prior knowledge of the target individual is available. VIPBench includes 22 specific target identities and a total of 80,080 images, covering both real and forged samples. These forgeries are generated using 14 state-of-the-art methods, spanning 7 face-swapping (FS) and 7 entire-face synthesis (EFS) techniques, providing diverse manipulation types and realistic evaluation settings. By centering evaluation around known identities and incorporating fine-grained annotations, VIPBench allows models to be assessed not only on detection accuracy but also on their ability to leverage identity-specific cues.
@@ 2.1 General Deepfake Detection
Current deepfake detection faces significant challenges in generalization. To illustrate, researchers have explored a range of approaches, including data augmentation [35, 33, 51, 4], frequency-based cues [45, 40], identity-aware learning [14, 24], disentanglement [71, 69], reconstruction [56], and custom network designs [11]. Data augmentation has proven especially effective for improving generalization—for example, FWA [35] simulates warping artifacts, Face X-ray and SBI [33, 51] target blending boundaries, and SLADD [4] uses adversarial examples to challenge models. Despite these advances, most traditional detectors still offer only binary outputs without human-understandable explanations, leaving users unclear about why a face is classified as fake—limiting trust and explainability.
@@
Vision and language are two core modalities in human perception, driving growing interest in visual-language multimodal learning. In deepfake detection, several studies [28, 50, 72] have explored prompt-based strategies for face forgery analysis, showing that multimodal large language models (MLLMs) offer better explainability than traditional detectors. Others [19, 32, 61, 30, 6, 74] have investigated different MLLMs for explainable detection, while [34] introduced a labeled multimodal dataset to support fine-tuning. However, most of these methods are designed for general-purpose detection and overlook the valuable identity-specific information available in many real-world scenarios.
@@ 2.2 Personalized Deepfake Detection for Specific Identity Protection
To detect identity inconsistencies in forged faces, prior works [49, 18, 16, 15, 9] use reference images for personalized deepfake detection. For example, ICT-Ref [15] employs a transformer to detect mismatches between inner- and outer-face regions, while DiffID [73] uses reconstruction-based identity distances to identify fakes. However, these methods mainly rely on global identity features and overlook deeper, user-specific information. This limits their robustness to distribution shifts and prevents them from offering human-understandable explanations.
@@ 3 VIPBench: A New Benchmark for Personalized Deepfake Detection
To promote the training and evaluation of personalized deepfake detection, we built a comprehensive identity-aware benchmark called VIPBench for personalized deepfake detection. The training set of VIPBench progressively fine-tunes MLLMs, advancing from basic facial attribute recognition to fine-grained identity inconsistency detection. Moreover, we introduce a new evaluation dataset [^2] We will release the dataset upon acceptance. for identity-aware deepfake detection, a setting that currently lacks sufficient evaluation resources, to assess the effectiveness of different methods. We obtained facial images from LAION-Face [75], CrossFaceID [59], and FaceID-6M [58], subjected to some preprocessing (details in supplementary materials). The main idea of the dataset construction pipeline is described below, while complete details are provided in the supplementary materials
Figure 3: Illustration of the proposed VIPBench, which includes three personalized datasets, (a) Facial Attributes Description Dataset $\mathcal{D}_{FA}$, (b) Identity Discrimination Dataset $\mathcal{D}_{ID}$, and (c) VIPEval $\mathcal{D}_{Eval}$. (d) Some examples of the facial attributes used in the $\mathcal{D}_{FA}$ are also illustrated here, while the full set is available in the supplementary material.
@@
By reformulating personalized deepfake detection as a target-face-centric, fine-grained face recognition problem, we constructed the Identity Discrimination Dataset $\mathcal{D}_{ID}$. This dataset, comprising $\mathcal{D}_{ID}^{general}$ and $\mathcal{D}_{ID}^{vip}$, includes facial images and corresponding annotations intended for reasoning about fine-grained identity discrimination. Specifically, $\mathcal{D}_{ID}$ comprises facial image pairs: positive (same-identity, real-real) and negative (all others). As shown in the ‘Data Collection’ part of Figure  2, we built facial pairs centered on VIP users in $D_{ID}^{vip}$, while facial pairs in $\mathcal{D}_{ID}^{{general}}$ were built using arbitrary identities. Given the high resemblance of forged faces to genuine ones, which is challenges for conventional face recognition, we augmented negative samples using SimSwap [5] (face swapping) and Arc2Face [44] (entire face synthesis). These negative pairs include different-identity real-real pairs and real-versus-forgery pairs, with these two categories balanced roughly 1:1. As demonstrated in Figure  3 (b), for each image pair, we first generated a facial attribute list using our fine-tuned Qwen2.5VL-7B model [1] $G$, trained on $\mathcal{D}_{FA}$. Then, we acquired the facial similarity scores by a pre-trained facial recognition model $f$. Gemini 2.5 Pro [^4] Gemini API version in use: 2.5-pro-exp-03-25. [20] was then strictly required to reason identity discrepancies between image pairs based solely on the provided facial attributes $G(I_{i})$ and $G(I_{j})$ and similarity scores $f(I_{i},I_{j})$, ensuring that the analysis was grounded in real features and not influenced by hallucinated content. The above process can be formulated as
$\begin{split}\text{Prompt}&=\text{Gemini}\left(Q,f(I_{i},I_{j}),I_{i},I_{j},G(I_{i}),G(I_{j})\right)\\ \mathcal{D}_{ID}&=\left\{I_{i},I_{j},\text{VQA}\left(I_{i},I_{j},\text{Prompt}\right)\;\middle|\;I_{i},I_{j}\in\mathcal{J}\right\},\end{split}$ (2)
@@
VIPEval: User-Specific Evaluation Dataset
We introduce a user-specific evaluation dataset called VIPEval for assessing user-specific forgery detection performance. Conventional datasets [36, 48] typically lack a sufficient number of high-resolution real images per individual, along with corresponding variations across different face forgery methods, which makes it challenging to evaluate personalized forgery detection approaches. The dataset includes images from rich sources and various resolutions with diverse generation methods, making it more representative of real-world conditions. To address this limitation, we carefully selected 22 unique identities from the previously described image-name pool $\mathcal{J}$, ensuring no overlap with the identities in the $\mathcal{D}_{ID}^{general}$. For each identity, 40-60 real images were collected, 20 reserved for testing in this benchmark $\mathcal{D}_{Eval}$, and the remainder used to construct $\mathcal{D}_{ID}^{vip}$. The benchmark includes reserved real images and a comprehensive set of forged counterparts. For each of the 22 test identities, 420 images are generated per method using seven distinct face-swapping (FS) techniques [5, 52, 27, 62, 64, 60, 17]. Additionally, three open-source entire-face synthesis (EFS) methods [44, 25, 22] generate 10 images per real test image by varying random seeds or prompts, resulting in 200 images per identity. Furthermore, four commercial API-based EFS methods [43, 29, 55, 31] produce 20 images per identity. In total, the dataset comprises 80,080 images.
@@ 4 VIPGuard: A Multimodal Framework for Personalized Deepfake Detection
4.1 Overview
We consider a scenario where only a few authentic images of a specific user (e.g., a celebrity, hereafter referred to as a VIP) are available. The objective is to model user-specific facial characteristics to identify suspicious images and protect the VIP’s identity from forgery attacks. To address the challenge, we propose a framework, VIPGuard, which develops an MLLM capable of identifying suspicious images of a specific user, while providing human-understandable explanations based on the user’s unique facial attributes. We reformulate personalized deepfake detection as a fine-grained face recognition problem centered on the protected target, where forgeries are detected through the MLLM’s reasoning over global identity features and detailed facial attributes. To equip the model with this capability, we start from a pre-trained MLLM (Qwen-2.5-VL-7B [1]) and progressively fine-tune it through a three-stage process: Face Attribute Learning, Identity Discrimination, and User-Specific Customization. We describe each stage in detail below.

### v2

@@ Abstract
Securing personal identity against deepfake attacks is increasingly critical in the digital age, especially for celebrities and political figures whose faces are easily accessible and frequently targeted. Most existing deepfake detection methods focus on general-purpose scenarios and often ignore the valuable prior knowledge of known facial identities, e.g., "VIP individuals" whose authentic facial data are already available. In this paper, we propose VIPGuard, a unified multimodal framework designed to capture fine-grained and comprehensive facial representations of a given identity, compare them against potentially fake or similar-looking faces, and reason over these comparisons to make accurate and explainable predictions. Specifically, our framework consists of three main stages. First, we fine-tune a multimodal large language model (MLLM) to learn detailed and structural facial attributes. Second, we perform identity-level discriminative learning to enable the model to distinguish subtle differences between highly similar faces, including real and fake variations. Finally, we introduce user-specific customization, where we model the unique characteristics of the target face identity and perform semantic reasoning via MLLM to enable personalized and explainable deepfake detection. Our framework shows clear advantages over previous detection works, where traditional detectors mainly rely on low-level visual cues and provide no human-understandable explanations, while other MLLM-based models often lack a detailed understanding of specific face identities. To facilitate the evaluation of our method, we build a comprehensive identity-aware benchmark called VIPBench for personalized deepfake detection, involving the latest 7 face-swapping and 7 entire face synthesis techniques for generation. Extensive experiments show that our model outperforms existing methods in both detection and explanation. The code is available at https://github.com/KQL11/VIPGuard .
@@ 1 Introduction
The rapid advancement of generative AI techniques [rombach2022high, tian2024emo, xu2024vasa, yan2025gpt, roop, deepfacelab] has led to the widespread creation and dissemination of deepfake content—synthetic media where a person’s identity is manipulated or replaced, often without consent. These techniques severely threaten personal reputation and public trust, especially for high-profile individuals such as celebrities, political figures, and public officials [bohacek2022protecting]. To this end, protecting one’s identity from such manipulations has become more than just a personal privacy issue—it is a pressing societal concern.
@@
Although deepfake detection has gained attention, most existing methods are designed for general-purpose use [qian2020thinking, li2020face, shiohara2022detecting, yan2023ucf, yan2024transcending, ma2025specificity, zhou2023exposing, zhou2025breaking, yu2023learning, yu2024diffforensics, yu2025reinforced]. They aim to classify any face image or video as real or fake, without considering whose face is being targeted. In the real world, many scenarios provide access to prior knowledge about the target identity [^1] In this work, we refer to such individuals as ”VIPs” whose identities we aim to protect., such as in the case of public figures or known individuals [ict_ref]. This opens a new avenue for detection: Can we leverage the known facial identity to improve both detection and explainability in personalized deepfake detection? Rather than treating all faces the same, identity-aware detection methods [id1, id2, id3, ict_ref, diff_id] focus on the semantic alignment between the input image and the authentic identity. By doing so, they provide more personalized and context-aware predictions, potentially improving both detection performance and explainability.
@@
However, existing identity-aware detectors [ict_ref, diff_id] fail to fully utilize the detailed identity-specific information. They primarily rely on global facial features while neglecting fine-grained semantic details—such as eye shape, facial contours, or other attributes. As shown in Figure 1, an image generated by the latest GPT-4o [gpt4o] can appear highly realistic at first glance, but still contains subtle inconsistencies in local facial regions—for example, unusually pronounced eye bags. When the target identity is known and all facial details are available, leveraging these fine-grained discrepancies for detection becomes especially promising.
@@
Figure 1: An illustrative comparison between natural images (from LAION-Face [laion_face]) and images generated by GPT-4o [gpt4o], showing localized inconsistencies in facial attributes, such as eye pouches and facial shapes.
@@
To this end, in this paper, we propose VIPGuard, a unified multimodal framework for detecting and explaining deepfakes targeting specific users. VIPGuard addresses identity-aware deepfake detection by explicitly incorporating known facial identity priors, including both global identity information and detailed structural facial attributes from the VIPs. To achieve this, this paper, for the first time, reformulates forgery detection as a fine-grained face recognition task, targeted at VIP identities. To leverage both global and local facial information, we use pre-trained face models [cosface, arcface, transface] to extract global facial priors (face similarity scores) and local facial priors (facial attributes). Leveraging these priors, we aim to enable an MLLM to perform forgery detection through semantic comparisons between suspect and authentic faces across visual and textual inputs. We train VIPGuard in three stages: (1) Fine-tuning an MLLM with a variety of facial attribute data to enhance facial understanding; (2) Performing discriminative learning to distinguish some arbitrary identities from manipulated or similar-looking faces by reasoning; (3) Supporting personalized detection by learning a unique and lightweight VIP token to represent each target identity for customized reasoning.
@@
To enable robust evaluation in personalized deepfake detection, we additionally present VIPBench—a comprehensive and identity-centric benchmark that differs fundamentally from conventional deepfake benchmarks [DeepfakeBench_YAN_NEURIPS2023, yan2024df40]. While existing benchmarks typically treat faces generically and overlook whose identity is being manipulated, VIPBench focuses explicitly on identity-aware scenarios, where prior knowledge of the target individual is available. VIPBench includes 22 specific target identities and a total of 80,080 images, covering both real and forged samples. These forgeries are generated using 14 state-of-the-art methods, spanning 7 face-swapping (FS) and 7 entire-face synthesis (EFS) techniques, providing diverse manipulation types and realistic evaluation settings. By centering evaluation around known identities and incorporating fine-grained annotations, VIPBench allows models to be assessed not only on detection accuracy but also on their ability to leverage identity-specific cues.
@@ 2.1 General Deepfake Detection
Current deepfake detection faces significant challenges in generalization. To illustrate, researchers have explored a range of approaches, including data augmentation [li2018exposing, li2020face, shiohara2022detecting, chen2022self, chen2025dual], frequency-based cues [qian2020thinking, luo2021generalizing, zhou2023exposing], identity-aware learning [dong2023implicit, huang2023implicit], disentanglement [yang2021learning, yan2023ucf], reconstruction [wang2021representative, yang2025all], and custom network designs [dang2020detection]. Data augmentation has proven especially effective for improving generalization—for example, FWA [li2018exposing] simulates warping artifacts, Face X-ray and SBI [li2020face, shiohara2022detecting] target blending boundaries, and SLADD [chen2022self] uses adversarial examples to challenge models. Despite these advances, most traditional detectors still offer only binary outputs without human-understandable explanations, leaving users unclear about why a face is classified as fake—limiting trust and explainability.
@@
Vision and language are two core modalities in human perception, driving growing interest in visual-language multimodal learning. In deepfake detection, several studies [jia2024can, shi2024shield, ye2024loki] have explored prompt-based strategies for face forgery analysis, showing that multimodal large language models (MLLMs) offer better explainability than traditional detectors. Others [foteinopoulou2025hitchhiker, xu2025fakeshield, xu2025avatarshield, li2024forgerygpt, wen2025spot, kang2025legion, zhang2024common, he2025vlforgery] have investigated different MLLMs for explainable detection, while [li2024fakebench] introduced a labeled multimodal dataset to support fine-tuning. Moreover, $\mathcal{X}^{2}$-DFD [chen2024x2dfd] further investigates hybrid frameworks that integrate conventional visual models with MLLMs. However, most of these methods are designed for general-purpose detection and overlook the valuable identity-specific information available in many real-world scenarios.
@@ 2.2 Personalized Deepfake Detection for Specific Identity Protection
To detect identity inconsistencies in forged faces, prior works [id1, id2, id3, ict_ref, cozzolino2021id, ming2025critical] use reference images for personalized deepfake detection. For example, ICT-Ref [ict_ref] employs a transformer to detect mismatches between inner- and outer-face regions, while DiffID [diff_id] uses reconstruction-based identity distances to identify fakes. However, these methods mainly rely on global identity features and overlook deeper, user-specific information. This limits their robustness to distribution shifts and prevents them from offering human-understandable explanations.
@@ 3 VIPBench: A New Benchmark for Personalized Deepfake Detection
To promote the training and evaluation of personalized deepfake detection, we build a comprehensive identity-aware benchmark called VIPBench for personalized deepfake detection. The training set of VIPBench progressively fine-tunes MLLMs, advancing from basic facial attribute recognition to fine-grained identity inconsistency detection. Moreover, we introduce a new evaluation dataset for identity-aware deepfake detection, a setting that currently lacks sufficient evaluation resources, to assess the effectiveness of different methods. We obtain all facial images from open-sourced datasets, including LAION-Face [laion_face], CrossFaceID [crossfaceid], and FaceID-6M [faceid6m], subjected to some preprocessing (details in supplementary materials). The main idea of the dataset construction pipeline is described below, while complete details are provided in the supplementary materials.
Figure 3: Illustration of the proposed VIPBench, which includes three personalized datasets, (a) Facial Attributes Description Dataset $\mathcal{D}_{FA}$, (b) Identity Discrimination Dataset $\mathcal{D}_{ID}$, and (c) VIPEval $\mathcal{D}_{Eval}$. (d) Some examples of the facial attributes used in the $\mathcal{D}_{FA}$ are also illustrated here, while the full set is available in the supplementary material. The real images shown in (c) are from CelebDF [laion_face] and VIPBench, while the fake ones are generated using multiple models. All images in (d) are sourced from LAION-Face [laion_face].
@@
By reformulating personalized deepfake detection as a target-face-centric and fine-grained face recognition problem, we constructed the Identity Discrimination Dataset $\mathcal{D}_{ID}$. This dataset, comprising $\mathcal{D}_{ID}^{general}$ and $\mathcal{D}_{ID}^{vip}$, includes facial images and corresponding annotations intended for reasoning about fine-grained identity discrimination. Specifically, $\mathcal{D}_{ID}$ comprises facial image pairs: positive (same-identity, real-real) and negative (all others). As shown in the ‘Data Collection’ part of Figure  2, we built facial pairs centered on VIP users in $D_{ID}^{vip}$, while facial pairs in $\mathcal{D}_{ID}^{{general}}$ were built using arbitrary identities. Given the high resemblance of forged faces to genuine ones, which is challenge for conventional face recognition, we augmented negative samples using SimSwap [simswap] (face swapping) and Arc2Face [arc2face] (entire face synthesis). These negative pairs include different-identity real-real pairs and real-versus-forgery pairs, with these two categories balanced roughly 1:1. As demonstrated in Figure  3 (b), for each image pair, we first generated a facial attribute list using our fine-tuned Qwen2.5VL-7B model [bai2025qwen2] $G$, trained on $\mathcal{D}_{FA}$. Then, we acquired the facial similarity scores by a pre-trained facial recognition model $f$. Finally, we utilized a commercial model (Gemini 2.5 Pro [^3] Gemini API version in use: 2.5-pro-exp-03-25. [Gemini] was selected in this paper), denoted as API in Figure  3 (b), to form the training data. Specifically, the API was then strictly required to reason identity discrepancies between image pairs based solely on the provided facial attributes $G(I_{i})$ and $G(I_{j})$ and similarity scores $f(I_{i},I_{j})$, ensuring that the analysis was grounded in real features and not influenced by hallucinated content. The above process can be formulated as
$\begin{split}\text{Prompt}&=\text{API}\left(Q,f(I_{i},I_{j}),I_{i},I_{j},G(I_{i}),G(I_{j})\right)\\ \mathcal{D}_{ID}&=\left\{I_{i},I_{j},\text{VQA}\left(I_{i},I_{j},\text{Prompt}\right)\;\middle|\;I_{i},I_{j}\in\mathcal{J}\right\},\end{split}$ (2)
@@
VIPEval (User-Specific Evaluation Dataset):
We introduce a user-specific evaluation dataset called VIPEval for assessing user-specific forgery detection performance. Conventional datasets [li2019celeb, rossler2019faceforensics++] typically lack a sufficient number of high-resolution real images per individual, along with corresponding variations across different face forgery methods, which makes it challenging to evaluate personalized forgery detection approaches. The dataset includes images from rich sources and various resolutions with diverse generation methods, making it more representative of real-world conditions. To address this limitation, we carefully selected 22 unique identities from the previously described image-name pool $\mathcal{J}$, ensuring no overlap with the identities in the $\mathcal{D}_{ID}^{general}$. For each identity, 40-60 real images were collected, 20 reserved for testing in this benchmark $\mathcal{D}_{Eval}$, and the remainder used to construct $\mathcal{D}_{ID}^{vip}$. The benchmark includes reserved real images and a comprehensive set of forged counterparts. For each of the 22 test identities, 420 images are generated per method using seven distinct face-swapping (FS) techniques [simswap, blendface, inswapper, uniface, xu2022mobilefaceswap, wang2021hififace, ghost3]. Additionally, three open-source entire-face synthesis (EFS) methods [arc2face, huang2024consistentid, guo2024pulid] generate 10 images per real test image by varying random seeds or prompts, resulting in 200 images per identity. Furthermore, four commercial API-based EFS methods [gpt4o, jimengai, TongYiWanXiang, kling] produce 20 images per identity. In total, the dataset comprises 80,080 images.
@@ 4 VIPGuard: A Multimodal Framework for Personalized Deepfake Detection
4.1 Problem Formulation and Comparison with Prior Works
We consider a personalized deepfake detection scenario in which the detector has access to several authentic images of the target user (e.g., a celebrity, hereafter referred to as a VIP), along with a collection of real facial images from other unrelated individuals. These authentic images serve as prior knowledge for the detector to better recognize and protect the target individual. The objective is to model user-specific facial characteristics to identify suspicious images and protect the VIP’s identity from forgery attacks. For forgeries, such as face-swapped images, we impose a realistic constraint that the source identities used for manipulation (denoted as $ID^{source}$) are unseen by the detector during training. This assumption is closer to the real-world scenario, where attackers can use arbitrary faces that are not available to the defender. This setting contrasts with identity-aware detection methods [ict_ref, id2, ming2025critical], which assume that both the target identity ($ID^{vip}$) and source identities ($ID^{source}$) are known during training—i.e., all relevant faces are included in the reference set $\{ID^{vip},ID^{source},ID^{others}\}$. Such a closed-world assumption simplifies the problem but is rarely realistic in practical applications. In our formulation, the detector only has access to $\{ID^{vip},ID^{others}\}$ during training, while $ID^{source}$ remains unknown. This difference introduces a more challenging yet practical problem setting, emphasizing the need for generalization to unseen source identities and better alignment with real-world deployment scenarios.
4.2 Method Overview
To address the challenge, we propose a framework, VIPGuard, which develops an MLLM capable of identifying suspicious images of a specific user, while providing human-understandable explanations based on the user’s unique facial attributes. We reformulate personalized deepfake detection as a fine-grained face recognition problem centered on the protected target, where forgeries are detected through the MLLM’s reasoning over global identity features and detailed facial attributes. To equip the model with this capability, we start from a pre-trained MLLM (Qwen-2.5-VL-7B [bai2025qwen2]) and progressively fine-tune it through a three-stage process: Face Attribute Learning, Identity Discrimination, and User-Specific Customization. We describe each stage in detail below.